Tharaldson Corporation makes a product with the following standard costs:

Standard Quantity or Hours Standard Price or Rate Standard Cost Per Unit
Direct materials 5.5 ounces $ 2.00 per ounce $ 11.00
Direct labor 0.4 hours $ 11.00 per hour $ 4.40
Variable overhead 0.4 hours $ 8.00 per hour $ 3.20

The company reported the following results concerning this product in June.

Originally budgeted output 4,100 units
Actual output 4,100 units
Raw materials used in production 20,500 ounces
Purchases of raw materials 21,600 ounces
Actual direct labor-hours 550 hours
Actual cost of raw materials purchases $ 42,800
Actual direct labor cost $ 14,100
Actual variable overhead cost $ 4,050

The company applies variable overhead on the basis of direct labor-hours. The direct materials purchases variance is computed when the materials are purchased.

The labor rate variance for June is:

(A) $1,523 F

(B) $8,050 F

(C) $1,523 U

(D) $8,050 U
Answer #1

Answer : D) $8,050 U

Labor Rate Variance = ( Actual Rate - Standard Rate )* Actual Hours

Actual  rate = 14,100/550 = 25.63

Labor Rate Variance = (Actual rate * actual hours)-( Standard Rate * Actual rate )

= 14,100-(11*550) = $8050 U